OpenStudy (anonymous):

first use appropriate properties of logarithms to rewrite and then find f(x)=x+5ln6x

OpenStudy (zzr0ck3r):

a*ln(b) = ln(b^a) so rewrite the expression and tell us what you got

OpenStudy (anonymous):

Here are some properties of logarithms: \[\large ln(x^n) = n\cdot ln(x)\] \[\large ln(x\cdot y) = ln(x) \cdot ln(y)\]

OpenStudy (zzr0ck3r):

I don't understand "and then find..." find what?

OpenStudy (zzr0ck3r):

@wio ln(x⋅y)=ln(x)⋅ln(y) is not true

OpenStudy (zzr0ck3r):

ln(a*b) = ln(a) + ln(b)

OpenStudy (anonymous):

find f'(x)

OpenStudy (anonymous):

oops \[ln(x\cdot y) = ln(x) + ln(y)\]

OpenStudy (zzr0ck3r):

correct, ok so do the simplifications we showed you and then show us what you got...

OpenStudy (anonymous):

Another one is: \[\Large n = ln(e^n)\]

OpenStudy (anonymous):

ln(5)+ln(6)?